A 65.0-kg skier is skiing down a 35.0o slope. If the friction force on her ski is 32.0N, what is her acceleration down the slope?

Answer:     m/s2

What is the normal force on the skier?
Answer #1

force acting down the plane = mg sin

= 65*9.8*sin 35 = 365.4 N

net force acting along the plane = (365.4 - 32)= 333.4 N

acceleration = net force / mass

= 333.4 / 65 = 5.13 m/s^2

Normal force = mg cos

= 65*9.8*cos 35 = 521.8 N